OpenStudy (anonymous):

Find the values of the trigonometric functions of theta from the given information tan(theta)=-4, sin(theta)>0

OpenStudy (dumbcow):

tan(theta) = -4 theta = arctan(-4) = 104.036 degrees

OpenStudy (anonymous):

I need to find the values such as the 6 trig functions sin, cos, csc, sec, cot

OpenStudy (icecube):

tan(theta) = -4 theta = arctan(-4) = pi - arctan(4) =104.036deg

OpenStudy (icecube):

sin(theta) = 4/sqrt(17) cos(theta) =-1/sqrt(17) cosec(theta) =sqrt(17) sec(theta) =-sqrt(17) cot(theta) = -1/4

OpenStudy (anonymous):

not quite... but close enough thanks!!